auto disiplin profit mahjong ways2 awsbet auto scatter merah mahjong ways auto tren harian scatter mahjong ways sss sistem rtp modern pgsoft sss strategi gila jam ajaib sss teknologi masa depan pgsoft sss terbukti akurat inilah pola dan jam sss terbukti valid begini cara cek rtp sss terkuak jam gacor mahjong selalu sss tersembunyi pgsoft bikin pola konsisten pola scatter hitam mahjong ways abc1131 rtp live mahjong ways abc1131 strategi mahjong wins 3 meledak petir merah scatter hitam mahjong ways auto fakta harian mahjong ways auto pusat kemenangan mahjong ways auto spam turbo spin awsbet ayu ilustrator digital buat temen tercengang kemenangan 99jt mahjong ways 1 qris deddy tukang masak viral kantongi rp81jt mahjong wins bermain situs magic134 gibran tukang fotografi menang rp344jt mahjong ways bermain situs hebitop hasil luar biasa bahlil raih 87jt kemenangan mahjong ways 2 lewat hebitop ss dari zero to hero ss fakta mengejutkan pola dan jam ss fakta tersembunyi dibalik rtp ss fitur auto maxwin ss fitur keren dan mode rahasia jackpot mahjong putaran awal mahjong gak perlu ditekan tekan mahjong cuan kilat tanpa modal besar pemulung anomali mahjong wins x10000 nelayan dermaga mahjong ways x1000 auto kode scatter merah mahjong ways auto mahjong wins3 awsbet cuan auto teori panen scatter mahjong ways mpoxo link slot mahjong mpo slot slot depo 5k slot deposit 1000 slot thailand asiaklub macauklub pondok88 garuda76 heylink macauklub heylink asiaklub heylink hksbet kapten76 heylink garuda76 heylink pondok88 heylink timpondok88 heylink mpoxo mpoxo AWSBET rawit303 mpoxl viral asiaklub viral macauklub viral garuda76 viral pondok88 magic134 ASLI777 Asli777 Mpogalaxy MPOGALAXY sakura76 baru slot gacor hari ini bathroomremodelingidea djakarta-miningclub abc1131 abc1131 ABC1131 abc1131 At-Taujih; Jurnal Bimbingan Konseling Islam MPO SLOT rtp mpoxo abc1131 slot viral abc1131 - sopadec abc1131 slot777 lundbergdesign.com

FMARS-15 XO Report 29-07-2023

 In XO Report

Sol: 6

Summary Title: Winterizing, Organizing, and Habitating

Author’s Name: D. Terry Trevino

Mission Status: All Nominal

Sol Activity Summary: Today’s activities were sporadic, but we kicked into gear, and activities picked up here on Devon Island, where we are settled and working away. Towards mid-day, the winds whipped up and are heavy, and the temperatures are creeping downward. The forecast appears to be very cold over the coming days, so we have started winterizing the research station as best we can.

EVA Report: Two relatively short EVAs were conducted today. The first led by Commander Andrew and Andy as EVA2 accompanied by bear guard Caleb. This EVA aimed to revisit a hydrothermal chimney identified back in 2017. The hydrothermal chimney was located and confirmed by Andrew who collected extensive photographs, a soil sample and a lichen sample. On the return the data loggers were photographed. Later in the day, EVA 6 was led by Olivia and Terry with Andrew as bear guard. An algal mate was revisited that was spotted on EVA4 and water, snow and algal samples were collected for cyanobacteria and microplastic studies. 

Back at the habitat, we awoke to fresh air; all air quality measurements were nominal and good. Andrew decided to test the results and closed up the station. After 10 hours, all the numbers began to elevate back to conditions requiring opening the station back up. The team is considering alternatives, including planning for the next analogue season a ventilation system or perhaps an ECLSS system.

Look Ahead Plan: As we proceed to Sol 7, our planned tasks include:

1. Continued Grounds Maintenance and winterizing of the station.

2. Flight Preparations for the Monday pickup of Andrew and Olivia: We will continue to properly pack the, trash-organize and weigh the expected loads for the plane flight scheduled on July 31. Furthermore, we will double-check the landing strip and send images of the strip to Borek Air.

3. EVA Preparation: we have a catch-up day for lost time and furthering our lab studies on Sunday. We will also ensure the facility is thoroughly cleaned.

4. Habitat Maintenance: Today, the installation of 6 LED light fixtures in the upper level of the hab on 2 independently switched runs was completed, providing controllable and cozy yet bright lighting for the common area. Full consumption appears to be approximately 60W, and switches allow cutting that in half if desired.

2 loads were incinerated thus far with 2 more planned this evening.

The windsock is performing well in sustained winds of 7m/s.

The wind turbine generator output is 18 volts and 2 amps.

Downstairs engineering area was further organized, with improvements including mounting of hook brackets to store extension cords.

Research and Reporting: Olivia and I continue our research on using fluorescent photography on the hunt for nano-plastics using the water samples on hand.

The required reports, including Sol Summary, Operations Report, Journalist Report, EVA Report, and EVA Request, will be diligently prepared and submitted as scheduled.

Weather: Mars on Devon Island experienced sunny and bright weather today, with temperatures around 10°C. The humidity remained low, and a brisk wind speed of 10 m/s during the mid-morning and throughout the day.

Crew Physical Status: The overall physical health of the crew remains nominal.

Support Requested: None requested.

Mission Success and Mars Exploration: The work continues as we explore the Simulated Red Planet, gathering samples and insights into methods for collecting these samples, which contribute to future missions and humanity’s long-term presence on Mars.